### v1

# A Non-Traditional Approach to Assisting Data Address Translation

Abstract
This paper proposes a novel way to assist conventional data address translation. The approach, PC-Indexed Data Address Translation (PCAX), uses the PC of a load instruction, and not a data virtual address, to obtain the page table entry (PTE) for the data accessed by a load instruction. PCAX is intended to be used for a small subset of the static loads in a program. We observe that: (i) a small subset of static loads is responsible for most of the misses in a data translation lookaside buffer (DTLB), and (ii) often a dynamic instance of a static load instruction accesses the same PTE as the last dynamic instance, and consider PCAX for this subset. With PCAX the effective miss rate of a conventional DTLB can be cut down by a factor of 2-3X in many cases, and even more in some cases. PCAX is also beneficial in reducing the number of secondary TLB (STLB) misses. Since the tables used for PCAX can be accessed alongside instruction fetch, they can be slow, yet frequently provide a valid PTE even before the data address calculation. This results in a performance improvement, and reduced data address translation energy, in most cases.
I Introduction
Translation from virtual to physical memory addresses is a fundamental operation that is carried out on every data memory access and is critical to the efficient processing of memory load (and store) instructions. A page table entry (PTE) in a page table holds a translation from a virtual page number (VPN) to a physical page number (PPN), along with other information. Data Translation Lookaside Buffers (DTLBs), accessed using the data address (virtual page number) of a load (or store) instruction, are used to cache PTEs to provide fast access and improve address translation efficiency. Given the criticality of data address translation to overall system performance, a plethora of techniques spanning both hardware and (system) software, have been proposed over the past several decades to improve its performance and efficiency.
Since the processing of load (and store) instructions is governed by the addresses of the data accessed, most techniques to improve their processing, including DTLBs for address translation, are based upon the empirical properties of the addresses of the data memory accesses. However, some techniques have been proposed to optimize load instruction processing that rely on the identity of the instruction—the instruction’s program counter (PC)—and not on the data memory addresses. A prime example is memory dependence prediction (MDP), also called memory disambiguation in some commercial products, which is routinely used in high-end processors to improve the scheduling of load instructions. In MDP, pairs of load and store instructions, identified by their PCs, are predicted to be dependent, agnostic of the actual data memory addresses that they access [31, 30, 14]. These pairs are tracked in a relatively small PC-indexed table which is consulted to predict whether a load could be beneficially executed before prior stores, without the determination of the data addresses of the load and store instructions.
In a similar vein, in this paper, we propose a non-traditional approach, PC-Indexed Data Address Translation (PCAX), to assist conventional data address translation which employs a conventional DTLB (which we refer to as a CDTLB in this paper). The approach relies on using the identity of an instruction (the instruction PC), and not on the actual data address to obtain a PTE for the data address that is expected to be accessed by the instruction. PCAX is intended to be used for a small subset of load instructions: those whose data translation access is expected to miss in a CDTLB.
PCAX is based on two important and novel empirical observations. First, a small percentage (and number) of static load instructions of a program account for a significant number of CDTLB misses. Second, quite often, when a dynamic instance of a static load has a CDTLB miss, it accesses the same virtual page (and the same PTE) as the previous dynamic instance. These observations suggest that a (relatively small) table of PTEs, accessed with the PC of a load instruction, a PC-Indexed Data Address Translation Table (PCAT), could provide the PTEs that would normally be obtained as a result of a DTLB miss. PCAX employs PCATs (loosely) coupled to the instruction caches. A PCAT is selectively accessed during instruction fetch and provides, with a reasonable accuracy, a needed PTE before the data address is generated, the CDTLB accessed, and a miss occurs. A PCAT is not meant to replace a CDTLB but rather augment it, by providing the PTE for a reference that likely will miss in the CDTLB.
Empirical results show that PCAX can be extremely effective, reducing the effective CDTLB miss rate by 2-3X in most cases for the benchmarks considered. PCAX is also useful in reducing the secondary TLB (STLB) misses. These miss reductions lead to improvements in both performance and energy efficiency in most cases.
II Motivation
A static load is a load at a given PC, a dynamic load is a dynamic instance of a given static load. A problem load is a static load whose dynamic instances result in CDTLB misses.
To motivate PC-indexed address translation, we present empirical observations on some characteristics of misses in conventional DTLBs (CDTLBs) when the memory address of the data for a load instruction is submitted to the CDTLB. For benchmarks, we consider a set of 84 traces from Qualcomm, for the ARM architecture. These traces, details of which are presented in Section IV, have been used in many recent papers [6, 12, 39, 5]. We consider a single page size of 4KB, and the miss rates, with a 64-entry CDTLB, range from 1% to 31%, with CDTLB MPKIs ranging from 1 to 53. (More on the traces in Section IV-A).
Figure 1: Percentage of static loads that are problem loads
Figure 1 shows the percentage of static loads that account for 90% of the: (i) accesses, and (ii) misses in a CDTLB. In the figure, the X-axis is the different benchmarks, and the Y-axis (note the log scale) is the percentage of static loads. We consider 3 CDTLB sizes: 64, 256 and 1024 entries, with 4-way associativity in each case. The larger sizes are representative of L2 or secondary TLBs. For a given CDTLB size, we only plot the cases for which the miss rate is greater than 1%.
The observation we make from the figure, which we do not believe has been documented before, is that a very small percentage of static loads are problem loads accounting for most (90%) of the misses in the CDTLB111 A related observation that has been known for a while is that a small percentage of static loads account for a large percentage of data cache misses [4, 51, 48]). This percentage of problem loads is typically an order of magnitude smaller than the percentage of static loads accounting for the references. For example, with a 64-entry CDTLB, in the median case, about 30% of the static loads account for 90% of the CDTLB accesses, but only about 4% of the static loads account for 90% of the CDTLB misses. From Figure 1 we also see that as the CDTLB size increases the percentage of static loads that are problem loads decreases  [^2] In rare cases the number of problem loads is higher for bigger CDTLBs. We address this in section IV-D. Overall, the percentage of problem loads is in the single digits.
Figure 2: Problem loads accessing same PTE
Figure 2 shows the percentage of time that a dynamic instance of a static load identified as a problem load accesses and misses on the same PTE as the last missing dynamic instance of the static (problem) load. The figure has 3 different DTLB sizes, 64, 256, and 1K entries. From the figure, we see that, for a very large percentage of the problem loads, the PTE that is accessed on a miss is the same as the PTE for the prior instance of the problem load. For example, at a median point, with a 64-entry CDTLB, about 75% of the loads that miss in the CDTLB access the same PTE as the prior dynamic instance. This percentage decreases as the CDTLB size increases. But even for the large CDTLB sizes, more than half of the problem loads access the same PTE as the prior dynamic instance.
The empirical observations of Figures 1 and 2 suggest that if we were to track the last PTE for static loads that miss in the CDTLB, in a table indexed by the load PC, we could get the correct PTE for the data that a load instruction would reference, and would likely miss in the CDTLB, a fairly large percentage of the time. Further, using the load PC, this correct (and CDTLB-missing) PTE could be obtained while the load was being fetched, i.e., even before it is decoded and well before the address of the data it will reference is determined. This tracking needs to be done only for a small percentage of the static loads in a program [^3] We will consider the absolute number, rather than the percentage of problem loads later when we evaluate different structure sizes..
III PC-Indexed Address Translation
III-A Overview
Motivated by the empirical observations of Section II, we propose PC-Indexed Address Translation (PCAX). The idea is to maintain PTEs for a select group of load instructions in a PC-Indexed Address Translation Table (PCAT). When one of these select loads is fetched, the PCAT is accessed to find a corresponding entry, the PTE from which is then provided in the expectation that it will be the PTE that will be needed when the address of the data referenced by that load is determined. The purpose of PCAX is to augment a CDTLB. For most loads, the CDTLB provides the PTE for the accessed data. For select load instructions, a (hopefully good) translation is provided by a PCAT.
III-B PCAT Structure and Operation
A PCAT is a cache, with a tag array and a data array. An entry shown in Figure 3, has (i) tag bits, (ii) a virtual page number (VPN), (iii) the PTE for that virtual page, which we collectively abbreviate as VPN-PTE, and (iv) two additional bits to help with replacement decisions: a Useful bit and a STLBM bit. Additionally, each entry has a Valid bit. The tag bits and the VPN-PTE are self-explanatory. We will see the use of the other bits later.
Figure 3: PCAT Tag and Data Entry
Since access to a PCAT entry is closely tied to an instruction access—a PCAT entry won’t be accessed if the corresponding load instruction isn’t also accessed—it is natural for the PCAT structures to mirror the cache structures holding instructions. Accordingly, we propose a 2-level PCAT hierarchy, an L1 PCAT (PCAT1) coupled to an L1 instruction cache and an L2 PCAT (PCAT2) coupled to an L2 cache.
Entries are created in the PCAT1 when there is a miss in the CDTLB. The PC of the load instruction is used to determine a place (set and way) where the new entry will reside, and the tag bits populated. The PTE returned on the CDTLB miss, and the VPN of the accessed data address, are entered into the VPN-PTE bits of the entry, and the Useful bit is unset. If the entry is created as the result of a Secondary TLB (STLB) miss, the STLBM bit is set; we discuss its use below. When an entry in the PCAT1 provides a useful translation, the Useful bit gets set.
A newly-created PCAT1 entry moves to PCAT2 when it is evicted from PCAT1 thus creating a new entry in the PCAT2, eventually populating the PCAT2  [^4] An alternative is to create an entry in the PCAT2 at the same time as in PCAT1. Eviction from the PCAT1 is independent of eviction of the corresponding code cache block from the L1i cache. However, an entry is moved from PCAT2 to PCAT1 (resulting in an eviction of an entry from PCAT1) when the corresponding instruction block is moved from the L2 to the L1i cache, and not on a PCAT1 miss, as in a typical cache hierarchy.
Figure 4: Relevant Portion of Load Operation
III-C Replacement in the PCATs
The PCATs can use normal cache replacement policies to determine a victim entry. However, we have found it beneficial to enhance a normal (e.g., LRU) replacement policy with some additional information. To preserve entries that have been deemed to be useful over those that have not been deemed so, we use the Useful bit to exclude an entry from a replacement decision when possible. Likewise, we have found it beneficial to try to preserve entries that were created as a result of an STLB miss, especially for the PCAT2, since they can potentially provide a translation that might otherwise have to be obtained via an STLB miss. The STLBM bit indicates such entries. The modified LRU replacement algorithm that we employ does not choose an entry with either its Useful or STLBM bit set, unless it doesn’t have a choice.
III-D Load Pipeline with PCATs
Figure 4 shows the steps with relevant portions for the operation of a load in the instruction pipeline, which we describe below. The PCAT1 is accessed using the PC of a load instruction, alongside the fetch of the load instruction. However, we only want to access the PCAT1 for loads for which there is an entry in the PCAT1, and not for every load instruction. The latter is wasteful as most load PCs won’t have a matching entry in the PCAT. For this, we can maintain a Problem Load (PL) bit with an instruction in the L1i cache. However, few instructions in the L1i cache are problem loads and thus few PL bits are set. Indeed, even for cache blocks that contain a problem load, often there is only one problem load. Accordingly, we use an Access PCAT (APT) bit with a block in the L1i cache. When an L1i cache block is accessed, and its APT bit set, the PCAT1 is accessed.
Though the PCAT1 is accessed alongside instruction fetch, its output—a VPN-PTE—isn’t needed until much later in the instruction processing pipeline. In particular, until the address for the data of the load instruction has been determined (and the VPN known). Accordingly, we propose the well-known two-step process for a cache access which is used to reduce access energy (at the expense of latency): access the tag array in the first step to determine if there is a hit, and the data array in the second step in case of a hit. On an access and a match, the PCAT1 provides a VPN-PTE, which is taken to the data address generation stage of the pipeline and maintained in a Recent PCAT Accesses Buffer (RPA buffer).
When the base register for a load instruction is read to start a data address calculation (or even earlier depending upon the pipeline) the VPN bits of the base register are compared with the VPN bits of the RPA buffer entries. On a match, the PTE from the RPA buffer is used if the VPN does not change in the address calculation (i.e., offset smaller than a page and no carry into VPN bits). In this way the access of the RPA buffer can be carried out in parallel with (or prior to) the address calculation. The VPN-PTE also contains the permission bits for the page (like a CDTLB) and permission checks are made at this point. If there is no match the data address is submitted to the CDTLB to obtain the corresponding PTE.
The VPN-PTE obtained from the PCAT1, while intended for use by the load instruction whose PC was used to access the PCAT1, can also be used by other instructions in the data address generation stage of the pipeline while it resides in the RPA buffer, i.e., until it is replaced by a new VPN-PTE from the PCAT1. As has been observed by others, different load instructions that are close in the instruction stream, and that use the same base register, often access the same virtual page (and PTE) [7, 13]. Similarly we have observed that different static loads that are temporally close in the dynamic instruction stream access the same PTE (they generally have the same base register, with different offsets, and likely access different fields of the same data structure). However, they are not all ”contiguous” in the dynamic stream of memory referencing instructions; they are interleaved with other memory referencing instructions which access other VPNs (that hit in the CDTLB). We have found this RPA buffer, which fleetingly holds select PTEs, to service the address translations for a fair number of instructions, as we shall see in section IV-I.
III-E Address bits for PCAT access
The bits used for a PCAT access could be from the PC of a (problem) load instruction, or the bits of the cache block number containing a (problem) load instruction. The former allows more flexibility as an arbitrary number of loads from the same cache block can reside in a PCAT. However, this requires a per-instruction APT bit in the L1i cache and also multiple accesses to the PCAT1 if these loads reside in different sets of the PCAT1. In the latter case, there needs to be only one APT bit, and one PCAT1 access per cache block. Multiple VPN-PTE entries for multiple problem loads in a cache block can reside in the different ways of a set and thus be obtained in parallel, with the set associativity capping the number of problem loads per cache block for which the PCAT1 can maintain information. We briefly evaluate the two options in section IV-E.
III-F PCAX Coupling with the CDTLB/STLB
A question that arises is whether the structures for PCAX (PCAT1/PCAT2) and the conventional address translation (CDTLB/STLB) should be coupled or not. That is, when a PTE is brought into the PCAT1, as a result of a CDTLB/STLB miss, should the PTE also be placed in the CDTLB/STLB (the structures are coupled), or solely in the PCAT1 (the structures are not coupled). Not coupling the structures frees up capacity in the CDTLB/STLB which could be used for other translations whereas not placing a PTE in the CDTLB when it is brought into the PCAT1 results in misses for other proximate load instructions that need the same PTE. Such instructions can’t get the translation from the PCAT1 (because of different PCs) and it won’t be there in the CDTLB.
Our experimental observations suggest that PCAT1/CDTLB coupling is beneficial for most applications as many other load instructions also access a PTE (from a CDTLB) brought in by a CDTLB-missing load. Accordingly, when a PTE is placed in the PCAT1, it is also placed in the CDTLB. For the STLB, given that there is little reason to maintain inclusion between the STLB and the CDTLB/PCATs, we do not couple the PCATs and the STLB. This reduces unnecessary duplication of entries in the PCAT1/CDTLB and STLB, and unnecessary evictions from the STLB, leading to a lower SLTB miss rate, as we shall quantify in section IV-H.
III-G PCAT/CDTLB/STLB and Page Table Interactions
Like a conventional CDTLB, a PCAT operates like a cache of PTEs, albeit one that is accessed with a load PC rather than a data address. When PTEs are cached, appropriate actions need to be taken to ensure correctness. When an update is made in a page table entry, copies cached in CDTLBs must be invalidated using a TLB Shootdown [47, 36]. These changes include the PPN mapping, permissions, and potential updates to the Access bit (if the page replacement algorithm requires it).
Since the PCAT is accessed with a load PC, and not with a data virtual address, operations that rely on a data virtual address (if contemplated) are more involved. TLB shootdown operation typically uses the data virtual addresses to perform invalidations. We invalidate all the entries in the PCATs on any change to an entry in the page table (a PCAT Shootdown). Since a PCAT entry is created on a (first) CDTLB miss, where the page walker also updates the Access bit in the page table and gives the correct value to the CDTLB, the PCAT will also have the correct value of the Access bit. Any subsequent clearing of the Access bit in the page table (by the OS) would result in a PCAT Shootdown if needed. Thus having a PCAT does not impact the operation of the page replacement algorithm.
The Dirty bit is more involved since the first write to a page can happen after an entry has been created in the PCAT. Moreover, the same PTE could be present with different load PCs in the PCATs. If a PTE is brought into the CDTLB from the PCAT (with an unset Dirty bit), rather than from the STLB or the page table (with a set Dirty bit), it simply results in an additional ”store miss” to set the Dirty bit in the CDTLB which then frequently hits in the STLB (with the Dirty bit set), only proceeding to the page table on an STLB miss.
III-H Periodically Invalidating the PCATs
The PCATs maintain VPN-PTEs associated with load instructions. The (valid) VPN-PTEs are always correct since any update to a PTE will result in the invalidation of the entire PCATs, as above. However, some performance (not correctness) pathologies can result because unproductive entries can be held on to in the PCATs longer than needed and reduce their effective size. These situations can arise when a load instruction with a PCAT entry becomes less active (or cold). One example we have observed is that when multiple entries for a code cache block are brought into the PCAT1 from the PCAT2, they include entries associated with cold load instructions which are not used in the PCAT1, but are considered to be accessed by the PCAT2 replacement algorithm, and thus not evicted from the PCAT2. Another example is when the number of entries per cache block is limited (e.g., by PCAT associativity), entries created by dynamically earlier instructions that become cold take up slots that later, more active instructions, could profitably use. To handle such situations created by dynamically cold code we can periodically clear the PCATs and re-populate them again. Such periodic invalidation and re-population have also been found to be useful in memory dependence predictors [14, 31], where store-load relationships that have been learnt are cleared to remove relationships that have become cold.
III-I Selectiveness of Loads Using PCAT
PCATs are intended to assist CDTLBs by providing translations that might miss in the CDTLB. PCAT sizes can be kept small because the number of static loads that miss in a CDTLB are small. However, PCAX could also be used for a small number of other ”high value” loads from the set of loads that would normally hit in the CDTLB (whose number is, of course, much larger). Like the data of figure 2 for loads that miss in a CDTLB, we have also found that many times a dynamic instance of a load that hits in a CDTLB uses the same PTE as the last dynamic instance, as we shall see in section IV-D. Judiciously using PCAX for loads that might normally hit in a CDTLB, while maintaining reasonably small-sized PCATs, to improve translation latency and energy is left for future work.
IV Evaluation
TABLE I: Simulated Machine Parameters
Processor Decoupled Front-end
Fetch Target queue 192 instr
Fetch Width 6 instructions
Decode queue 60 entry
Dispatch queue 60 entry
Branch target buffer 8K entries
Target cache 4K entries
Return Address stack 64 entries
Branch penalty 2 cycles (decode stage)
Branch Predictor Hashed perceptron
Processor Back-end
Execute width 4 instructions
Retire width 5 instructions
Re-order buffer 352 entries
Load, store queue 128, 72 entries
Memory hierarchy
CPU Frequency 4GHz
L1 ITLB 64-entry, 4-way, 1-cycle, 4-entry MSHR
L1 DTLB 64-entry, 4-way, 1-cycle, 4-entry MSHR
L2 TLB 1536-entry, 12-way, 8-cycle;
4-entry MSHR, 1 page walk / cycle
Page Structure Caches 3-level Split PSC, 2-cycle; PML4: 2-entry, fully;
PDP: 4-entry, fully; PD: 32-entry, 4-way
L1 Dcache 48KB, 12-way, 5-cycle, 8-entry MSHR;
next line prefetcher
L2 Cache 512KB, 8-way, 10-cycle, 16-entry MSHR;
spp-dev prefetcher
LLC 2MB, 16-way, 20-cycle, 32-entry MSHR
DRAM 4GB, one 8-byte channel, 1600 MT/s;
tRP=tRCD=tCAS=12.5
IV-A Benchmarks and Experimental Infrastructure
For our evaluation, we use ChampSim [18], a detailed simulator that simulates an out-of-order core with 6-wide fetch and 4-wide issue. Table I summarizes the parameters of the simulated microarchitecture. This version of ChampSim also models a realistic page table walker as found in x86 architectures. Key aspects modeled include: (i) variable latency cost of page table walks and (ii) caching of page walk references in data memory hierarchy as in [50]. Specifically, this models a 4-level page table, a hardware page table walker, and a 3-level split page structured cache. The page walker supports up to 4 concurrent STLB misses. All caches and TLBs use a normal LRU replacement policy. The default page size is 4KB and the default PCAX scheme employs a 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA Buffer. PCAT1 uses normal LRU while PCAT2 uses the slightly modified LRU replacement which was explained in Section III-C.
For benchmark programs, we use a set of 84 server traces from Qualcomm, which were released after modifications recently [17, 16]. These traces have been widely used in recent papers using the Champsim simulation infrastructure, and studying a variety of issues in instruction supply, memory hierarchies, and microarchitecture [50, 49, 6, 39, 28, 23] and more which can be found in [20]. Though the data footprint of these traces—in the range of 20MB-100MB—isn’t as large as others that have been used in studies of data address translation [26, 25, 21], the data referencing patterns result in a range of data TLB miss rates, from 1% to 31% with a 64-entry, 4-way set associative, data TLB, making them an apt choice for evaluating PCAX. All traces are run to completion with a warm-up of 10M instructions. Many benchmarks traces have 100M instructions, while some of the traces are smaller (30M instructions) [3].
IV-B Storage Costs
An entry in the PCAT1 takes 11.6B (20-bit tag, 1 valid bit, 1 useful bit, 1 STLBM bit, 38-bit virtual page, and 4B for the PTE). A 64-entry PCAT1 takes about 744B. PCAT2 with 512, 1K, and 2K entries takes about 6KB, 12KB, and 24KB, respectively. There are 512 APT bits in the 32KB, 64B block, L1i cache, taking up 64B. There are 8192 APT bits in the 512KB, 64B block, L2 cache, taking up 1024B. The RPA buffer takes up 18B.
We use a configuration with a 64 entry PCAT1 and 2K PCAT2 which amounts to a total storage of 25.8KB. However, even 512 and 1K entry PCAT2 provide similar performance for many of the applications, as we show in Section IV-F.
IV-C Results Overview
We start our evaluation by presenting some relevant characteristics of the benchmark programs in section IV-D. Then section IV-F evaluates the reduction in miss rates for primary DTLBs for different PCAT configurations. Section IV-G studies the effect on larger page sizes. Section IV-H evaluates the reduction in miss rates for secondary DTLBs for different PCAT configurations. Section IV-I presents the service sites for data address translation for a select benchmark set and then sections IV-J and IV-K evaluate the impact on CPI and energy, respectively. Lastly, Section IV-L studies the effect of invalidating the PCATs periodically.
The ordering of the benchmarks in the different figures is different for better visual presentation.
Figure 5: Total Data Pages Touched
Figure 6: CDTLB Load Miss Rate
Figure 7: Number of Problem Loads
Figure 8: Maximum Problem Loads in L1i
IV-D Benchmark Characteristics
In this section, we consider some key benchmark characteristics that are relevant to our evaluation. Figure 5 presents the number of 4KB data pages touched by the different applications (note log scale on Y-axis). The data footprint is not very large. However, the benchmarks also have large code footprints [6, 5, 12, 39, 38]; for an L1i of size 32KB, there are 45 benchmarks with an L1i MPKI between 0-30, 11 benchmarks with an L1i MPKI between 30-45, 21 benchmarks with an L1i MPKI between 45-60 and 7 benchmarks with an L1i MPKI $>$ 60. The large code footprint results in very frequent code movement from L2 to L1i, and coupled with the data referencing pattern collectively results in a significant number of CDTLB misses, shown in Figure 6 for CDTLB sizes of 64, 128, 256, and 1024 entries, with 4-way associativity in each case. For a 64-entry CDTLB, most of the applications have a miss rate greater than 5%. Even for a (single-level) 1024-entry CDTLB, several applications have a miss rate greater than 2%.
Next, we consider problem loads that account for 90% of the CDTLB misses. Figure 7 shows the total number of problem loads for 64 and 128 entry CDTLBs. The total number is of the order of a few hundred and less than 2000 in almost all cases  [^5] In some cases the number is larger for a 128-entry CDTLB than for a 64-entry one. This is because those applications have a high miss rate with a 64-entry CDTLB (e.g., 20%), which comes down quite significantly (e.g., to 10%) with a 128-entry CDTLB. Accounting for 90% of the misses bring in more problem PCs, which aren’t included in the 90% when the miss rate is higher.. These numbers in this figure give an idea of how large of a PCAT2 might work well.
Since the PCAT1 needs to accommodate only problem loads that are resident in the L1 instruction cache, we consider this number in Figure 8. The figure presents the maximum number of problem loads resident in the 32KB L1 instruction cache at any time for 64-entry and 128-entry CDTLBs. This maximum number is in the small hundreds. The average number (not shown) is smaller. As we shall see, PCAT1 sizes that can accommodate a few tens of problem loads (e.g., 64) are adequate in most cases.
Figure 9: Loads Accessing same PTE
Figure 10: PTE Reuse Characteristics
Figure 11: PTE Rereference Distance
In Section II, we analyzed the load references that miss in the CDTLB on the same PTE as the previous dynamic missing instance. Figure 9 presents the percentage of all loads (which includes both hits and misses in the CDLTB) that access the same PTE as the previous dynamic instance of the load. We observe that this percentage is as high as 70-90% for many of the benchmarks. This result should not be surprising as many dynamic instances of a static load access proximate elements of a data structure, which are likely to be in proximate memory addresses, and thus mostly on the same page.
Figure 10 presents the reuse characteristics of PTEs, that is, after an initial load reference that brings a PTE into the CDTLB, what percentage of reuses of the PTE (either from another dynamic instance of the same static load, or from dynamic instances of other static loads) occur with $<$ 10, $<$ 50, $>$ 100, or $>$ 1000 dynamic load instructions. From the figure we see that a very large percentage of references (and reuses) to a PTE occur within fewer than 50 dynamic load references and a very small percentage are for $>$ 100 references. Further, we have observed, data not shown, that dynamic instances of an average of 2-8 other static loads reference a PTE brought into a CDTLB by a dynamic instance of a static load. These results should also not be surprising because we expect references to different fields of an element of a data structure (from different static loads), which are likely to be on the same page, to occur close to each other.
Figure 11 presents the re-reference distance for a PTE. That is, between the time a PTE is brought into the CDTLB by a dynamic instance of a static load, how many other PTEs are referenced, on average, until the time the same PTE is re-referenced and brought into the CDTLB again by another dynamic instance of the same static load. Data is presented for re-reference distances of greater than 100, 200, 500, and 1000 PTEs. As we see, many benchmarks have more than 50-60% of such missing PTE references with a re-reference distance greater than 100 and 200. Even the percentage of such misses with re-reference distances greater than 500 and 1000 is non-trivial (20-30%) for some of the benchmarks. Such large re-reference distances make it difficult to maintain a given PTE in a relatively small-sized CDTLB, leading to high miss rates.
To put it together, a dynamic instance of a static load references a given PTE, which is then (mostly) quickly followed by dynamic instances of other static loads referencing the same PTE. The load then gets replaced from the L1i and the PTE from the CDTLB (at different points of time). Later the same load is brought into the L1i again, and then it references the same PTE again. This is what provides the basis for PCAX: associate PTEs with a small subset of loads (which miss in the CDTLB) and bring a PTE associated with a load into the PCAT1 when the load is brought back into the L1i so that it can be used when the load references the PTE without having to incur a CDTLB miss. In other words, view the issue of CDTLB misses not from the viewpoint of the data reference stream, but from the viewpoint of the instruction reference stream which creates the data references that result in the CDTLB misses. If a PTE is associated with a load, it can be accessed with the load PC, not with the address of the data being accessed.
IV-E PCAT Indexing
Figure 12: PCAT Indexing
Figure 12 presents the CDTLB miss rates for cache block address and load PC indexing of a PCAT. The miss rates are very similar in most cases, with either performing slightly better in some cases. We use the block address indexing for all the results we present subsequently because it results in only a single PCAT access for a block of instructions.
IV-F Reduction in Primary CDTLB Misses
Figure 13: Effective Load Miss Rate (64-Entry CDTLB)
Figure 14: Effective Load Miss Rate (128-Entry CDTLB)
Figure 15: Comparing 128-entry CDTLB with PCAX
Figure 16: Varying PCAT2 size
We now consider the reduction in the primary CDTLB misses due to the translation requests fulfilled by a PCAT1. Figures 13 and 14 show the effective CDTLB miss rate for 64- and 128-entry, 4-way set associative CDTLBs, respectively. The effective miss rate is the miss rate for load references that can’t be serviced by the PCAT1 and CDTLB, and end up accessing the STLB. In each case, we plot the base CDTLB miss rate (Base) and the effective miss rate with 32-, 64-, 128-entry, 2-way set associative PCAT1s. The PCAT2 size is 2048 entries and is 8-way set associative in all cases.
The figures show a significant reduction in the effective CDTLB miss rate in almost all cases, even with a 32-entry PCAT1. A 64-entry PCAT1 can achieve more reduction in several cases; reduction beyond 64-entries is marginal. The magnitude of the reduction is quite impressive. From figure 13 observe that while the base miss rate is 4%-12% for most of the applications, it is brought down to 2%-5% with a 64-entry PCAT1. For many benchmarks, there is a 3X reduction in the effective miss rate. The significant improvements continue even for a 128-entry CDTLB in figure 14. There is a greater than 2X reduction in the miss rate with a PCAT, with base miss rates in the 4%-6% range brought down to the 2%-2.5% range for many applications.
Figure 15 combines select information presented in Figures 13 and 14 to allow for a direct comparison between a 64-entry CDTLB, a 128-entry CDTLB, and a 64-entry CDTLB and 64-entry PCAT1 (PCAX). The latter two take up essentially the same amount of storage. As we can see, PCAX outperforms the 128-entry CDTLB in reducing the miss rate in almost all cases. Later in sections IV-J and IV-K we will also see that PCAX gets better performance and consumes less energy, respectively, than the larger 128-entry CDTLB.
Next, we consider the impact of PCAT2 size in Figure 16. Here we have a 64-entry, 4-way set associative CDTLB and a 64-entry, 2-way set associative PCAT1. We have plots for the effective CDTLB load miss rate for PCAT2 configurations of 512, 1K, 2K, and 4K, with 8-way set associativity in all cases, in addition to the base CDTLB (Base) and also simulate a configuration with no PCAT2. PCAT2 sizes of 512 and 1K entries do quite well for many applications, but leave some opportunity for improvement in others. A 2K entry PCAT2 is nearly as effective as the ones with 4K entries. Referring back to Figure 7, barring conflicts due to limited associativity, given that most applications had fewer than 2000 problem loads, we expected a 2K entry PCAT2 to be quite effective. For many applications that had fewer than 500 problem loads, a 512-entry PCAT2 would be adequate. The results of Figure 16 bear this out. Not using a PCAT2 does not achieve the benefits of PCAX in several cases. This is because the PCAT entry is lost when the load is evicted from the L1i and thus is not available for reuse when the load is brought back into the L1i for processing.
IV-G Large Page Sizes
Figure 17: Large Page Study
Figure 18: Large Page Improvements with PCAX
Figure 19: Large Page Comparison with 4K+PCAX
We now consider the effectiveness of PCAX with (uniform) page sizes larger than 4KB. We consider only a uniform page size as our traces do not have multiple page sizes. Figure 17 presents the percentage of loads that miss in a 64-entry, 4-way CDTLB that accesses the same PTE as the previous dynamic instance. We consider page sizes of 4KB, 16KB, 64KB, 256KB, and 1MB, and present data for benchmarks where the CDTLB miss rate is greater than 1%. Clearly, the phenomenon that underlies PCAX—that a large number of misses access the same PTE as the previous dynamic instance—is prevalent even for larger pages. Figure 18 presents the miss rate improvement for larger page sizes using PCAX. The CDTLB miss rate improvements continue even for larger page sizes.
Figure 19 presents CDTLB miss rates for large pages and also for 4KB pages with PCAX. Clearly, using larger page sizes brings down the miss rate. More interestingly using 4K pages with PCAX provides lower miss rates compared to 16K and 64K page sizes. It performs as well for many benchmarks or comes close to the performance using a 256K page size. This suggests that with PCAX larger page sizes may not be needed if a reduction in CDTLB misses is the objective.
Figure 20: Effective STLB Miss Rate
Figure 21: Data Address Translation Service Sites
IV-H Reduction in STLB Misses
Because the PCATs are not impacted by the normal operation and replacements in the CDTLB and STLB, they are able to hold on to address translations that would miss in them for a longer period. Further, by not directly placing a PTE in the STLB on an STLB miss, and placing it in the PCAT1 and CDTLB instead, allows other (potentially useful) PTEs to continue to reside in the STLB, rather than be evicted. This helps reduce the number of STLB misses. Figure 20 shows the effective STLB load (global) miss rate for the base case, and for PCAT2 sizes of 512, 1K, 2K, 4K, and 8K entries, with 8-way set associativity in each case. The PCAT1 is a 64-entry, 2-way set associative. To improve readability, we only plot the applications where the Base miss rate is greater than 1%. With PCATs, we can achieve a noticeable reduction in the effective STLB (global) miss rate. Though the absolute magnitude is small, the percentage reduction is meaningful, especially as a miss in the STLB results in expensive page table accesses (Page Walks), which can involve several memory references.
IV-I Data Address Translation Service Sites
We now consider the different sites from where a data address translation is serviced in Figure 21. The applications in the figure represent a subset of the applications merely to improve readability, as presenting data for all 100 applications would be cluttered. Applications not considered show similar results, and we will consider all the applications when we factor this data into the CPI impact in section IV-J and energy impact in section IV-K. The results are for a 64-entry, 4-way CDTLB, 1536-entry, 12-way STLB, 64-entry, 2-way PCAT1, a 2K entry, 8-way PCAT2, and a 2-entry RPA buffer.
With conventional address translation, a valid PTE can be obtained from (i) the CDTLB (labeled CDTLB), (ii) the STLB in case of a CDTLB miss and STLB hit (labeled STLB), or (iii) the Page Table in case of an STLB miss (labeled Page Walk). The left bar for a benchmark in the figure shows the percentage of load references that get a PTE from these different sites. With a PCAT, a valid PTE could additionally be obtained from the PCAT in three different situations: (a) the load instruction that was used to access the PCAT1 and the translation access would have missed in the CDTLB (labeled PCAT-M), (b) the same load instruction and the translation access would have hit in the CDTLB (labeled PCAT-HS), or (c) a different load instruction whose translation access would have hit in the CDTLB (labeled PCAT-HD). For (b) and (c) the PTE is obtained from the RPA buffer. The right bar for a benchmark shows the percent of load instructions that get their address translation in each of the six different ways. The Y-axis uses a log scale so that the difference in the important portions, PCAT-M and Page Walk, can be better appreciated.
Observe that the number of references served by the CDTLB and the STLB are cut down significantly in most cases because of the references served by the PCAT (PCAT-M, PCAT-HS, and PCAT-HD). There is even a non-trivial reduction in expensive Page Walks in cases where STLB misses are non-negligible. These reductions in CDTLB and STLB accesses, and Page Walks, lead to a lower data translation CPI and energy as we quantify next.
IV-J CPI for Data Address Translation
In this section, we consider the improvement in the CPI for load data address translation with PCAX, borrowing from the CPI analysis in [26], and the potential performance benefit. The parameters of the CPI model are shown in Table II. Our computation of cycles spent in load address translation in conventional components consists of: (i) Cycles ($Cycles_{L1-H}$) spent in servicing CDTLB Hits ($H_{L1-H}$), where hit latency is 1 cycle (ii) Cycles ($Cycles_{L1-M}$) spent servicing CDTLB Misses ($M_{L1-M}$), which lookup the STLB that takes 8 cycles (iii) Cycles ($Cycles_{L2-M}$) spent servicing STLB misses ($M_{L2-M}$) which takes an average of 30 cycles. Since PCAX provides a PTE before the address calculation, we consider it to take 0 cycles for address translation.
Figure 22 shows the data address translation CPI for load instructions for a 64-entry, 4-way CDTLB (Base-64) and a 128-entry, 4-way CDTLB (Base-128) and a 64-entry, 4-way CDTLB with PCAX with default PCAT1/PCAT2 sizes (PCAX) for the benchmarks. PCAX achieves a lower effective CPI, with the 64-entry CDTLB and 64-entry PCAT1 configuration having a lower CPI than even a 128-entry CDTLB in most cases.
The lower data translation CPI also results in improved overall performance. Figure 23 shows the percentage improvement in performance (over Base-64), obtained from a timing simulation that incorporates a detailed TLB hierarchy and page walk modeling (as was used in [50, 49])666 The ordering of the benchmarks on the X-axis in Figure 23 is different from that in Figure 22.. The performance improvement is on par with many other microarchitectural techniques. For most benchmarks (except 5 of 84), the 64-entry CDTLB with PCAX configuration has better performance improvement than a 128-entry CDTLB (even though we optimistically model the same access latency for both CDTLB sizes).
TABLE II: Load Data Address Translation CPI Model
Performance Model
CDTLB Hits CyclesL1-H = 1*HL1-H
CDTLB Misses CyclesL1-M = 8*ML1-M
STLB Misses CyclesL2-M = 30*ML2-M
Total cycles CyclesTLB = CyclesL1-H+CyclesL1-M+CyclesL2-M
M: Misses H: Hits
Figure 22: CPI per load with PCAX
Figure 23: Performance improvement with PCAX
TABLE III: Dynamic Energy for Read/Write Operation
Component
Size
Assoc
Read
Write
CDTLB 64 4-way 0.79 1.11
CDTLB-128 128 4-way 0.91 1.72
STLB 1536 12-way 3.56 3.9
PCAT1 64 2-way 0.25 0.74
PCAT2 2048 8-way 1.3 3.09
L1-Cache 32KB 8-way 9.3 14.31
MMU-Cache-PD 32 4-way 0.69 0.9
MMU-Cache-PDP 4 Full 0.08 0.02
MMU-Cache-PML4 2 Full 0.04 0.02
Figure 24: Energy improvement with PCAX
IV-K Energy Impact
Next, we consider the energy impact of PCAX over a base 64-entry, 4-way CDTLB. We measure energy consumption using CACTI 7.0 [29] with 22nm technology. The dynamic energy per Read and Write Operation for the conventional address translation structures and the PCATs are shown in Table III. It is worth noting that the energies of the PCAT1 and PCAT2 are smaller than the CDTLB and STLB because they can be run a lot slower. In the case of PCAT1, it can take multiple cycles (5 in our case) to respond and still provide a translation before the address calculation. Likewise, PCAT2 (10 cycles) is accessed alongside the (slower) L2 cache. The longer latencies also allow for sequential tag and data access for energy savings, which we use for PCAT1 and PCAT2.
Figure 24 shows the impact of PCAX on the dynamic energy for data address translation 777 The ordering of the benchmarks on the X-axis in Figure 24 is different from the ordering in Figures 22 and 23.. The numbers in the figure account for both loads and stores. If we considered the energy for only loads, the energy benefit would be higher as stores have little energy benefit. The figure shows a small energy cost in a few benchmarks. These are also the cases that have a smaller number of CDTLB and STLB misses, to begin with, and PCAX provides a smaller reduction in the absolute number of misses. However, for most benchmarks, there is a meaningful improvement in dynamic energy, in addition to the CPI and performance benefits we saw earlier. The energy reduction is higher in cases when page walks are reduced due to fewer STLB misses, resulting in fewer L1D cache accesses due to fewer page walks. A 128-entry, 4-way CDTLB (Base-128), on the other hand, takes more energy than a 64-entry CDTLB (improvement is negative) in most cases.
IV-L Effect of Invalidating PCATs
Figure 25: Impact of Periodic PCAT Invalidation
Figure 25 studies the impact of periodically invalidating the PCATs. Invalidating the PCATs after 5-10M instructions has almost no effect, and invalidating every 1M instruction has a marginal impact in some cases.
V Related Work
Given the importance of data address translation, there has been a plethora of work in improving the efficiency of the process. The overarching theme of the vast body of work is to decrease the number of DTLB “misses”, i.e., the number of times an address translation requires accessing the secondary TLBs or the page tables. We provide a brief overview of some of this prior work below; excellent summaries of such work can also be found in some recent papers [21, 50, 22, 26].
There is a large body of work on novel DTLB designs. Some of these designs [37, 35, 45] seek to exploit contiguity in both the virtual and physical pages and minimize the number of TLB entries, thereby reducing the DTLB misses. Processors and the OS provide support [2] for large pages to increase the TLB reach and reduce the number of DTLB misses. Different TLB designs have been explored to better support huge pages [15, 43, 33, 26]. Huge pages rely on memory contiguity, which might be absent due to fragmentation and a plethora of techniques have been proposed to work around this concern [9, 34, 25, 19, 32, 44, 46]. Huge pages exploit discrete levels of contiguity, [22] is a scheme that allows exploiting intermediate levels of contiguity. Techniques to reduce TLB misses have also employed prefetching (primarily into the secondary TLB), exploiting memory access stream patterns [11, 24, 42]. A paper [50] proposes combining prefetchers and picking one of multiple depending on their usefulness and exploits locality in the page table to cache nearby PTEs and minimize page walks (and effectively prefetching PTEs). Another recent paper to increase TLB reach involves reorganizing the page table and using hash functions to significantly reduce miss rates for data-intensive applications [21] .
Techniques have also been proposed to reduce the TLB miss cost (mainly page walk cost) by caching parts of the page walk in the memory hierarchy. Such techniques employ MMU caches to cache higher levels of the page table and reduce the latency of a page walk [8, 10, 40, 1], and are found in many modern processors. PTEs are also cached to reduce page walk latency.
The VAX processors used the concept of pre-translation, which observed that multiple dynamic instances of instruction fetching accessed the same PTE and optimized it by maintaining the PTE in a special register [41]. Regarding optimizations to the processing of load instructions based upon their identity rather than the data address, researchers have also made an empirical observation that a small number of static load instructions account for a disproportionate number of data cache misses [4, 51, 48] and proposed appropriate optimizations. Instruction PC-based prediction has also been proposed to optimize coherence events due to store and load instructions in a multiprocessor [27].
VI Concluding Remarks
We presented a novel technique that uses the PC of a load instruction, rather than the address of the referenced data, to obtain an PTE for the referenced data. The key observation driving the use of a PC rather than the data address is that often a dynamic instance of a load instruction accesses the same virtual page as the prior dynamic instance, thus needing the same PTE. This technique can only be profitably used by a small subset of the static loads in the program. We further observed that only a small subset of the static loads accounts for most of the address translation misses with conventional DTLBs. Combining these two observations we proposed PC-Indexed Data Address Translation (PCAX), to assist conventional address translation structures. PCAX employs tables that require additional storage, which is modest given the transistor budgets in contemporary high-performance chips, but is able to achieve a significant reduction (2-3X or more) in the effective DTLB miss rate, and improve performance, while also reducing the energy for data address translation in most of the considered benchmarks.
A plethora of approaches have been proposed to reduce the DTLB miss rate, with varying degrees of effectiveness and with varying consequences for system software over the past several decades. PCAX is a purely microarchitectural solution that is highly effective, with no demands on the ISA, software and/or systems.
To limit the number of loads for which PCAX is applied, and keep the associated tables (reasonably) small, we limited the application of PCAX to loads that are likely to miss in a DTLB. It can also be applied to a few “high value” loads that would normally hit in a DTLB and thus further reduce the overall latency and energy for data address translation. Using the PC of an instruction to obtain a translation for the data that it will access also opens up the possibility of many other innovations for data address translation. We leave these to future work.